OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. m = 2, A(_, 0), B(5, 10) need the answer -5 0 5

OpenStudy (fortisdea13):

slope m = (10-0)/(5-x) = 2 10 = 2(5-x) 10 = 10-2x -2x = 10-10 = 0 x = 0 m = 2, A(0, 0), B(5, 10)

rebeccaxhawaii (rebeccaxhawaii):

so do you know the slope formula?